Use la place transform to solve Y-5y-by = U (+-1) given IVP Y(0)=0 Y(O) = 1
Answer #1

y s4 -by - ult-1 you-o, y 100= Apply captace on both sidos L194 sy-by g = (ult-11) sys- syloi - yto) - 51 syls - 403) - bylssos if i ffistle = f(t) then zor es FISI S (5²-56-6] I f e as fisil = (t al(t-a) , a=1 S152_566) If stress-b)} Apply partial14** ***** T = 116-1) (7 + 2 e (t-1) + 1 1 2 0 0 Apply partial fractions Sb 5² 55.6 [stills 6) (studis-6) 1 = R0 (5-6) + QxfS